A call and put expire in 0.41 year and have an exercise price of $100. The underlying stock is worth $90 and has a standard deviation of 0.25. The annual risk-free rate is 11 percent. The annual dividend yield (q) on the stock is 2%. The put option price from the three-period binomial model is:
Answer #1
As per Black Scholes Model
Value of put option = N(-d2)*K*e^(-r*t)-(S*e^(q*t))*N(-d1)
Where
S = Current price = 90
t = time to expiry = 0.41
K = Strike price = 100
r = Risk free rate = 11.0%
q = Dividend Yield = 2%
σ = Std dev = 25%
d1 = (ln(S/K)+(r-q+σ^2/2)*t)/(σ*t^(1/2)
d1 = (ln(90/100)+(0.11-0.02+0.25^2/2)*0.41)/(0.25*0.41^(1/2))
d1 = -0.34763
d2 = d1-σ*t^(1/2)
d2 =-0.34763-0.25*0.41^(1/2)
d2 = -0.507708
N(-d1) = Cumulative standard normal dist. of -d1
N(-d1) =0.635941
N(-d2) = Cumulative standard normal dist. of -d2
N(-d2) =0.694171
Value of put= 0.694171*100*e^(-0.11*0.41)-90*e^(-0.02*0.41)*0.635941
Value of put= 9.59